Suppose that the production function is given by 𝑦=0.5√𝐾√𝐿

a) Derive the steady-state levels of output per worker and capital per worker in terms of the saving rate, s, and the depreciation rate, δ.

b) Derive the equation for steady-state output per worker and steady-state consumption per worker in terms of s and δ.


1
Expert's answer
2021-11-22T09:57:32-0500

a)Given  Y=0.5√𝐾√𝐿

"\therefore \frac{y}{L}=\frac{0.5\sqrt K\sqrt L}{L}=\frac{0.5\sqrt K}{\sqrt L}"


"y=\frac{Y}{L}=0.5\sqrt \frac{K}{L}"


where y=output per worker and K=capital per worker.


Steady state occurs when "\Delta K=0=sy-\delta K"


"0.5s\sqrt K=\delta K=K=\frac{0.5s}{\delta}"


"K=\frac{0.25s^2}{\delta}" ..............steady state level of capital per worker.


"y=0.5\sqrt K=\frac{0.25s}{\delta}".................. steady state level of output per worker


b)

"y=c+i" where c=consumption per worker


i=investment per worker="sy=\frac{0.25s^2}{\delta}"


"c=y-i=\frac{0.25s}{\delta}-\frac{0.25s^2}{\delta}"


"c=\frac{0.25s}{\delta}(1-s)"..................... steady state level of consumption per worker